Find the perimeter of the square in which a circle of area 81 pi is inscribed.
How can I find that? Please show me.

Answer by user_dude2008(1862)

A=81pi ----> 81pi=pi*r^2 ----> 81=r^2 ----> r=9

side length of square = 2*r = 2*9 = 18

Perimeter = 4*18 = 72

Answer: 72